### v1

# On the single versus the repetitive Penrose process in a Kerr black hole

Abstract
Extracting the rotational energy from a Kerr black hole (BH) is one of the crucial topics in relativistic astrophysics. Here, we give special attention to the Penrose ballistic process based on the fission of a massive particle $\mu_{0}$ into two particles $\mu_{1}$ and $\mu_{2}$, occurring in the ergosphere of a Kerr BH. Bardeen et al. indicated that for the process to occur, some additional ”hydrodynamical forces or superstrong radiation reactions” were needed. This idea was further expanded by Wald and Chandrasekhar. This animosity convinced T. Piran and collaborators to move from a simple three-body system characterizing the original Penrose process to a many-body system. This many-body approach was further largely expanded by others, introducing additional processes, some questionable in their validity. In this letter, we return to the simplest original Penrose process and show that the solution of the equations of motion, imposing the turning point condition on their trajectories, leads to the rotational energy extraction from the BH expected by Penrose. The efficiency of energy extraction by a single process is precisely quantified for three different single decay processes occurring respectively at $r=1.2M$, $r=1.5M$, and $r=1.9M$. An interesting repetitive model has been proposed by Misner, Thorne & Wheeler [1, hereafter MTW73]. Indeed, it would appear that a repetitive sequence of $246$ decays of the above injection process at $r=1.2M$ and the corresponding ones at $r=1.5M$ and $r=1.9M$ could extract $100\%$ of the rotational energy of the BH. The accompanying paper shows that accounting for the irreducible mass introduces a non-linear approach that avoids violating energy conservation, leading to new energy extraction processes and demonstrating the impossibility of extracting the whole BH rotational energy by a sequence of Penrose processes.
preprint: APS/123-QED
I Introduction
The physics of a Kerr Black Hole (BH) is based on three contributions:
  1. 1.
    The mathematical solution of the Einstein equations by Roy Kerr [2] that plays a central role in the description of rotating objects in General Relativity.
  2. 2.
    The discovery by Carter of the separability of the Hamilton-Jacobi equations, describing geodesics in the Kerr spacetime, in terms of four conserved quantities [3, 4].
  3. 3.
    The introduction by Ruffini and Wheeler of the effective potential governing the motion of test particles in the BH equatorial plane [4, 5]
    $\frac{V}{\mu}=\frac{2\hat{a}\hat{p}_{\phi}\pm\sqrt{\hat{r}\left[\hat{a}^{2}+\hat{r}\left(\hat{r}-2\right)\right][\hat{p}_{\phi}^{2}\hat{r}+\hat{r}^{3}+\hat{a}^{2}(\hat{r}+2)]}}{\hat{r}^{3}+\hat{a}^{2}\left(\hat{r}+2\right)}.$ (1)
We use geometrical units throughout and consider the positive sign solution of Eq. (1). For the BH, we indicate by $M$ and $L$ the mass and angular momentum and by $\hat{a}=L/M^{2}$ the dimensionless spin parameter. Similarly, for a massive test particle of mass $\mu$ and angular momentum $p_{\phi}$ and energy $E$, we indicate the dimensionless quantities $\hat{E}=E/\mu$, $\hat{p}_{\phi}=p_{\phi}/(M\mu)$, and for the dimensionless radial coordinate is given by $\hat{r}=r/M$.
A particle with energy $E$ and $E>V$ is endowed with a positive kinetic energy with $\left(\frac{dr}{ds}\right)^{2}>0$. The particles with $E=V$ correspond to the turning point with $\frac{dr}{ds}=0$. For stable circular orbits, the conditions $V^{\prime}=0$ and $V^{\prime\prime}>0$ are satisfied, where a prime denotes derivate with respect to the radial coordinate (see Landau and Lifshits [4], Ruffini and Wheeler [5]).
From the above equations, it was possible to define the new concept of the ergosphere (see, e.g., Landau and Lifshits [4], Ruffini and Wheeler [5, 6]), which extends from the horizon
$\hat{r}_{+}=1+\sqrt{1-\hat{a}^{2}}$ (2)
to the stationary limit
$\hat{r}_{\rm ergo}=1+\sqrt{1-\hat{a}^{2}\cos^{2}\theta}\ .$ (3)
In 1969, Penrose proposed [7] the occurrence of a ballistic decay process between the horizon and the stationary limit of a Kerr black hole (BH) to extract its rotational energy. In this approach, a massive particle $\mu_{0}$ with energy $E_{0}$ decays into two particles $\mu_{1}$ and $\mu_{2}$. The particle $\mu_{1}$, with negative energy $E_{1}$ and negative angular momentum $p_{\phi 1}$, enters into the BH horizon, while the particle $\mu_{2}$ reaches back a distance at infinity with total energy $E_{2}$ larger than $E_{0}$. The difference of energy between $E_{2}$ and $E_{0}$ corresponds to the extraction of the rotational energy of the BH.
The first example of a ballistic process was in Fig. 2 of Christodoulou’s paper [8] (submitted on September 17, 1970). This figure summarizes the computation performed by Ruffini and Wheeler in [5], where some kinematic parameters were determined.
This result was possible by assuming that the decay process occurs at the three turning points of the particles’ trajectories. The equations from which this example was derived and the conditions of the turning point were later published [5]. Three months after the publication of Christodoulou’s paper and the appearance of Fig. 2, the Penrose-Floyd paper appeared [9] (submitted on December 16, 1970). The same equations were integrated, missing the crucial simplification of using the turning point condition. Therefore, only a general statement was inferred “The particular example described here is somewhat inefficient in this respect. ”.
Following this, a large number of papers appeared. The most quoted has been the one by Bardeen et al. Instead of integrating the numerical equations of the potential as seen from infinity, they went through a detailed analysis of the description of the trajectories by using an orthonormal tetrad formalism around every spacetime event. However, this work missed verifying the actual trajectories. They concluded that for the Penrose process to be viable, the outgoing particle should have a velocity $v_{2}>c/2$. Indeed, as we will see in Sec. III, the simplest ballistic Penrose process can easily fulfill this condition. Therefore, the above inequality does not constrain the validity process.
Instead, Bardeen et al. stated that ”… energy extraction cannot be achieved unless hydrodynamical forces or superstrong radiation reactions can accelerate fragments to more than this speed during the infall. On dimensional grounds, such boosts seem to be excluded …”. Chandrasekhar made an analogous statement in [10], page 371: ”In other words, the fragments must have relativistic energies before any extraction of energy by a Penrose process becomes possible”.
This led Piran et al. [11, 12, 13] to consider the many-particle Penrose collisional process. This multi-particle approach was correctly performed in their papers [11, 12, 13]. Many publications followed, extending further the multi-particle approach and additional considerations which were, however, not necessarily correct (see, e.g., [14] and [11]). As this article explicitly shows, these generalized Penrose processes are not necessary to prove the Penrose process feasibility.
In Sec. II, for general convenience, we formulate the laws of energy and angular momentum conservation, assuming the essential simplifying condition that the decay process for the three particles occurs at the turning points of the effective potential [8], simplifying the Penrose-Floyd approach.
In Sec. III, we perform the integration of the equations considering the specific example of an incoming particle $\mu_{0}$ with mass $10^{-2}M$ and performing the decay process very close to the horizon at $\hat{r}=1.2$. These choices respond to two opposite requirements: to test the most energetic process by performing the fission process by choosing a decay close to the horizon and also to fulfill the necessary conditions of the test particle approximation for the incoming particle, see Misner, Thorne & Wheeler [1, hereafter MTW73].
We determine the masses $\mu_{1}$ and $\mu_{2}$ implied by the conservation laws in Sec. II.
We focus on the change of the mass and angular momentum of the BH due to the capture of the particle $\mu_{1}$ by the BH and especially on the relativistic nature of the particle $\mu_{2}$, which is indeed ultra-relativistic and carries away a portion of the BH energy and angular momentum, so fulfilling the Penrose expectation and verifying the results. We also verify the decay process inside the ergosphere in three locations, i.e., in addition to $1.2M$, at $1.5M$ and $1.9M$. The results are summarized in Table 1 and Figs. 1, 2, and 3.
In Sec. IV, following an interesting suggestion by [MTW73], we perform an iterative Penrose process and verify that in all three above cases it is possible to extract the entire rotational energy of a Kerr BH in a finite number of iterative steps. The results are represented in Fig. 4.
However, to perform this process, the number of injected particles should be much larger than the BH mass and be unaccountable in the overall mass energy of the process. We are then confronted with a basic difficulty in the theoretical validity of the framework of the Penrose process itself. This problem is addressed in an adjoin publication in which the necessary implementation of the BH mass-energy formula is implemented [8, 15, 16], taking into account the irreducible mass $M_{\rm irr}$ and its relation to the horizon surface area
$S=16\pi M_{\rm irr}^{2}.$ (4)
This will lead to a highly nonlinear energy extraction process.
II The basic equations for the Penrose process
We return to the basic equations governing the physical process of the decay of an incoming massive particle “0”, with initial kinetic energy zero at infinite distance, into two massive particles “1” and “2” with masses
$\mu_{0},\ \ \mu_{1},\ \ \mu_{2}:\ \mu_{0}>\mu_{1}+\mu_{2}.$ (5)
Following [MTW73], we assume that the test-particle approximation holds with the mass ratio of the test particle being $\mu_{0}/M\leq 10^{-2}$ and, consequently, specifying the details of the energy extraction process.
We assume that the decay process occurs at the turning points in the ergosphere of an extreme Kerr BH, where particle “1” counter-rotates with the BH and is in a negative energy state, as measured from an infinite distance [4, 5, 6, 8, 17, 10]. Meanwhile, particle “2” returns to infinity with mass energy of $\mu_{2}\hat{E}_{2}$, which is larger than that of the incoming particle, $\mu_{0}\hat{E}_{0}$, potentially resulting in the extraction of rotational energy from the BH.
For simplicity, we consider the energy extraction from an extreme Kerr BH, $a/M=1$. The effective potential can be deduced from Eq. (1) as
$\frac{V}{\mu}=\frac{2\hat{p}_{\phi}+\left(\hat{r}-1\right)\sqrt{\hat{r}\left(\hat{r}^{3}+\hat{r}+2\right)+\hat{p}_{\phi}^{2}\hat{r}^{2}}}{\hat{r}^{3}+\hat{r}+2},$ (6)
which for $r/M\to 1$, $V/\mu\to\hat{p}_{\phi}/2$. For particle motion outside the horizon $\hat{r}>\hat{r}_{+}=1$, the effective potential gives an important property
$\hat{E}>\frac{\hat{p}_{\phi}}{2}.$ (7)
Which will be used in the following analysis.
From the basic equations for the extreme Kerr BH, by setting values of $\hat{r}$, $\hat{E}_{1}$ and $\hat{p}_{\phi 1}$ that fulfill $21<\hat{r}<$ 2, $\hat{E}_{1}<0$, $\hat{p}_{\phi 1}<0$, $\hat{E}_{1}>\hat{p}_{\phi 1}/2$, and masses that satisfy the mass condition (5), one can always find the corresponding positive values of geodesic conserved quantities $\hat{E}_{0}$, $\hat{p}_{\phi 0}$, $\hat{E}_{2}$, and $\hat{p}_{\phi 2}$ from the set of equations [18]
$\mu_{0}\hat{E}_{0}$ $=\mu_{1}\hat{E}_{1}+\mu_{2}\hat{E}_{2},$ (8a)
$\mu_{0}\hat{p}_{\phi 0}$ $=\mu_{1}\hat{p}_{\phi 1}+\mu_{2}\hat{p}_{\phi 2},$ (8b)
$\mu_{0}\hat{p}_{r0}$ $=\mu_{1}\hat{p}_{r1}+\mu_{2}\hat{p}_{r2},$ (8c)
where
$\hat{p}_{r0}$ $=-\left[(\hat{E}_{0}^{2}-1)\hat{r}^{3}+2\hat{r}^{2}\right.$
$+\left.(\hat{E}_{0}^{2}-\hat{p}_{\phi 0}^{2}-1)\hat{r}+2(\hat{E}_{0}-\hat{p}_{\phi 0})^{2}\right]^{1/2}=0,$ (9a)
$\hat{p}_{r1}$ $=-\left[(\hat{E}_{1}^{2}-1)\hat{r}^{3}+2\hat{r}^{2}\right.$
$+\left.(\hat{E}_{1}^{2}-\hat{p}_{\phi 1}^{2}-1)\hat{r}+2(\hat{E}_{1}-\hat{p}_{\phi 1})^{2}\right]^{1/2},$ (9b)
$\hat{p}_{r2}$ $=\left[(\hat{E}_{2}^{2}-1)\hat{r}^{3}+2\hat{r}^{2}\right.$
$+\left.(\hat{E}_{2}^{2}-\hat{p}_{\phi 2}^{2}-1)\hat{r}+2(\hat{E}_{2}-\hat{p}_{\phi 2})^{2}\right]^{1/2},$ (9c)
where Eqs. (8a)–(8c) represent the conservation of energy, angular momentum, and radial momentum, respectively. Equation (9a) guarantees that the particle “0” decays at the turning point. Notice that the above system of algebraic equations (8a)-(9c) is under-determined. Indeed, for a set of 12 unknowns, represented by the masses of the test particles as well as their specific energies, angular momenta, and radial momenta, there are only 7 equations. However, we are not free to select 5 unknowns arbitrarily since the chosen values must satisfy the boundary conditions of the process. We address only the maximum energy extraction process in agreement with the test-particle approximation.
Table 1: Change of the parameters of an extreme Kerr BH with mass $M$ and angular momentum $L=M^{2}$ due to the capture of a particle with negative energy $E_{1}=\hat{E}_{1}\mu_{1}=\Delta M$ and negative angular momentum $\Delta L=\hat{p}_{\phi 1}\mu_{1}M$. We list the results for the decay occurring at three selected positions inside the ergosphere: $\hat{r}=1.2$, $1.5$, and $1.9$. We choose $\hat{E}_{0}=1$, $\hat{p}_{\phi 1}=-19.434$, $\mu_{2}/\mu_{1}=0.78345$, and from the conservation laws, we obtain the other parameters. For $\hat{r}=1.2$: $\mu_{1}=0.0209\mu_{0}$, $\mu_{2}=0.01637\mu_{0}$, $\hat{p}_{\phi 0}=2.113$, $\hat{E}_{1}=-6.936$, $\hat{E}_{2}=69.942$, and $\hat{p}_{\phi 2}=153.872$. For $\hat{r}=1.5$: $\mu_{1}=0.0218\mu_{0}$, $\mu_{2}=0.0171\mu_{0}$, $\hat{p}_{\phi 0}=2.268$, $\hat{E}_{1}=-3.520$, $\hat{E}_{2}=62.995$, and $\hat{p}_{\phi 2}=157.485$. For $\hat{r}=1.9$: $\mu_{1}=0.0247\mu_{0}$, $\mu_{2}=0.0194\mu_{0}$, $\hat{p}_{\phi 0}=2.456$, $\hat{E}_{1}=-0.501$, $\hat{E}_{2}=52.326$, and $\hat{p}_{\phi 2}=151.737$. For all cases, we adopt $\mu_{0}=10^{-2}M$.
$\hat{r}$ $\frac{E_{2}}{M}$ $\gamma_{2}$ $\frac{E_{\rm extracted}}{M}$ $\frac{E_{1}}{M}=\displaystyle\frac{\Delta M}{M}$ $\frac{\Delta L}{M^{2}}$ $\frac{L_{f}}{M^{2}}$ $\frac{M_{f}}{M}$
$1.2$ $1.145\times 10^{-2}$ $69.942$ $1.449\times 10^{-3}$ $-1.449\times 10^{-3}$ $-4.061\times 10^{-3}$ $0.9959$ $0.9985$
$1.5$ $1.076\times 10^{-2}$ $62.99$ $7.681\times 10^{-4}$ $-7.681\times 10^{-4}$ $-4.240\times 10^{-3}$ $0.9957$ $0.9992$
$1.9$ $1.012\times 10^{-2}$ $52.326$ $1.236\times 10^{-4}$ $-1.236\times 10^{-4}$ $-4.799\times 10^{-3}$ $0.9952$ $0.9998$
Figure 1: Example of a particle decay into two particles in the ergosphere of an extreme Kerr BH occurring at $\hat{r}=1.2$. The three particles fulfill the condition (5) for $\mu_{1}/\mu_{0}=0.0209$ and $\mu_{2}/\mu_{0}=0.01637$. For these masses, and further assuming $\hat{E}_{0}=1.0$ and $\hat{p}_{\phi 1}=-19.434$, the solution of the conservation equations given in Sec. II leads to $\hat{p}_{\phi 0}=2.113$, $\hat{E}_{1}=-6.936$, $\hat{E}_{2}=69.942$, and $\hat{p}_{\phi 2}=153.872$. The upper right panel shows the particle “0” coming in from infinity. The left panel displays particle “1” with negative energy (represented by a black point) falling towards the horizon. The lower right panel shows the particle “2” returning to infinity after decay.
The specific energy of particle “0” ($\hat{E}_{0}$) can be equal to, larger, or smaller than $1$. We choose the parameters so particle “0” drops from rest at infinity. To use the general formula and make a specific example, we must determine the radius at which the decay happens within the ergosphere. The set of trajectories is determined for each value of $\hat{r}$.
When particle “1”, with negative energy and negative angular momentum, falls into the BH, it reduces the mass and angular momentum of the BH according by
$\Delta M=\hat{E}_{1}\mu_{1}$ (10)
and
$\Delta L=\hat{p}_{\phi 1}M\mu_{1},$ (11)
respectively. The energy extracted from the BH by particle “2” can be expressed as
$E_{\rm extracted}=\hat{E}_{2}\mu_{2}-\hat{E}_{0}\mu_{0}=-\hat{E}_{1}\mu_{1}$ (12)
We also note that $E_{\rm extracted}$ is maximal when particle “1” is created at the turning point near the horizon of the BH ($\hat{r}=1$) where $\hat{E}_{1}=\hat{p}_{\phi 1}/2$.
III The case of $\mu_{0}=10^{-2}M$ and consequences on the energy extraction process
We start assuming the dimensionless diagram of Fig. 1, in which a particle of negative energy $E_{1}$ and negative angular momentum $p_{\phi 1}$ is absorbed by an extreme Kerr BH with mass $M$ and angular momentum $L=M^{2}$.
Using the energy and angular momentum conservation laws as well as the condition that the decay occurs at the turning point of the three trajectories, i.e., $p_{r0}=p_{r1}=p_{r2}=0$, we find that the mass of the incoming particle $\mu_{0}$ can be left as an independent parameter. In contrast, the remaining masses of the test particles are derived through the conservation laws (8)–(9).
A new simplifying analytic approach avoids some numerical integration procedures. Assuming the incoming particle starts at rest from infinity, i.e., $\hat{E}_{0}=1$, that the decay occurs at some $\hat{r}$, and we supply a value of $\hat{p}_{\phi 1}$ and the outgoing to infalling particle mass-ratio $\mu_{2}/\mu_{1}$, the system of equations can be solved analytically for the rest of variables:
$\hat{p}_{\phi 0}$ $=\frac{2-(\hat{r}-1)\sqrt{2\hat{r}}}{2-\hat{r}},$ (13a)
$\hat{E}_{1}$ $=\frac{2\,\,\hat{p}_{\phi 1}+\sqrt{\rho+\hat{r}^{2}\Delta\,\hat{p}_{\phi 1}^{2}}}{R},$ (13b)
$\hat{E}_{2}$ $=\frac{\mu_{0}}{\mu_{2}}-\frac{\mu_{1}}{\mu_{2}}\hat{E}_{1},$ (13c)
$\hat{p}_{\phi 2}$ $=\frac{\mu_{0}}{\mu_{2}}\hat{p}_{\phi 0}-\frac{\mu_{1}}{\mu_{2}}\hat{p}_{\phi 1},$ (13d)
$\mu_{1}$ $=\mu_{0}\frac{\rho}{K+\sqrt{K^{2}-\rho^{2}[1-(\mu_{2}/\mu_{1})^{2}]}},$ (13e)
where
$\Delta$ $=(\hat{r}-1)^{2},\quad R=\hat{r}^{3}+\hat{r}+2,\quad\rho=\hat{r}\,\Delta\,R$ (14a)
$K$ $=\sqrt{(\rho+\hat{r}^{2}\Delta\hat{p}_{\phi 0}^{2})(\rho+\hat{r}^{2}\Delta\hat{p}_{\phi 1}^{2})}-\hat{r}^{2}\Delta\hat{p}_{\phi 0}\hat{p}_{\phi 1}.$ (14b)
Figure 2: Example of a particle decay into two particles in the ergosphere of an extreme Kerr BH occurring at $\hat{r}=1.5$. For $\hat{E}_{0}=1.0$, $\hat{p}_{\phi 1}=-19.434$, and $\mu_{2}/\mu_{1}=0.78345$, the solution of the conservation equations given in Sec. II leads to $\mu_{1}/\mu_{0}=0.0218$, $\mu_{2}/\mu_{0}=0.0171$, $\hat{p}_{\phi 0}=2.268$, $\hat{E}_{1}=-3.520$, $\hat{E}_{2}=62.995$, and $\hat{p}_{\phi 2}=157.485$. The upper right panel shows the particle “0” coming in from infinity. The left panel displays particle “1” with negative energy (represented by a black point) falling towards the horizon. The lower right panel shows the particle “2” returning to infinity after decay.
Table 1 and Figs. 13 summarize the results obtained from the solution given by Eqs. (13) and (14), assuming $\hat{p}_{\phi 1}=-19.434$ and $\mu_{2}/\mu_{1}=0.78345$, for three selected decay points, $\hat{r}=1.2$, $1.5$, and $1.9$. We can use these values to determine the parameters of the new BH, i.e., the final mass $M_{f}=M_{i}+E_{1}=M+\hat{E}_{1}\mu_{1}=0.9985M$ and the final angular momentum $L_{f}=L_{i}+L_{1}=M^{2}+\hat{p}_{\phi 1}\mu_{1}M=0.9959M^{2}$. Finally, the extracted energy becomes $E_{\rm extracted}=E_{2}-E_{0}=\hat{E}_{2}\mu_{2}-\mu_{0}=1.45\times 10^{-3}M$.
Notice that the energy of the incoming and outgoing particles can be represented as $E_{0}=\gamma_{0}\mu_{0}$ and $E_{2}=\gamma_{2}\mu_{2}$, respectively. Then, the corresponding dimensionless coefficients $\hat{E}_{0}$ and $\hat{E}_{2}$ can be identified with the Lorentz factor $\gamma_{0}$ and $\gamma_{2}$, respectively. It then follows that $\gamma_{0}=1$, indicating that the particle “0” has zero kinetic energy at infinity. For the particle “2” carrying out the rotational energy, we obtain $\gamma_{2}=69.942$, which corresponds to a velocity of approximately $99.989\%$ of the speed of light, i.e., the outgoing particle is ultra-relativistic. This means that, in this particular case, the outgoing particle acquires a clear signature with an enormous kinetic energy due to its interaction with the gravitational field of the BH. It is easy to verify that this result satisfies the inequalities established by Wald [19], and Bardeen et al. [17] (see also [10]).
Figure 3: Example of a particle decay into two particles in the ergosphere of an extreme Kerr BH occurring at $\hat{r}=1.9$. For $\hat{E}_{0}=1.0$, $\hat{p}_{\phi 1}=-19.434$, and $\mu_{2}/\mu_{1}=0.78345$, the solution of the conservation equations given in Sec. II leads to $\mu_{1}/\mu_{0}=0.0247$, $\mu_{2}/\mu_{0}=0.0194$, $\hat{p}_{\phi 0}=2.456$, $\hat{E}_{1}=-0.501$, $\hat{E}_{2}=52.326$, and $\hat{p}_{\phi 2}=151.737$. The upper right panel shows the particle “0” coming in from infinity. The left panel displays particle “1” with negative energy (represented by a black point) falling towards the horizon. The lower right panel shows the particle “2” returning to infinity after decay.
In Fig. 2, we illustrate the case in which an incoming particle with energy $E_{0}=\mu_{0}=10^{-2}M$ decays at a radius $\hat{r}=1.5$. By setting the same values of $\hat{E}_{0}$, $\hat{p}_{\phi 1}$ and $\mu_{2}/\mu_{1}$ of the previous example, we obtain in this case $\mu_{1}/\mu_{0}=0.0218$, $\mu_{2}/\mu_{0}=0.0171$, $\hat{p}_{\phi 0}=2.268$, $\hat{E}_{1}=-3.520$, $\hat{p}_{\phi 2}=157.485$ $\hat{E}_{2}=62.995$. A summary of the main quantities for this decay is given in Table 1.
Fig. 3 illustrates the case of an incoming particle that decays at $\hat{r}=1.9$. For the same assumptions, we obtain $\mu_{1}/\mu_{0}=0.0247$, $\mu_{2}/\mu_{0}=0.0194$, $\hat{p}_{\phi 0}=2.456$, $\hat{E}_{1}=-0.501$, $\hat{E}_{2}=52.326$, and $\hat{p}_{\phi 2}=151.737$. We refer to Table 1 for a summary of the properties of this decay.
The Penrose process remains the only consistent theoretical model to extract the rotational energy of a Kerr BH using a test particle of mass in the BH ergosphere with an efficiency larger than nuclear or chemical energy processes [20]. Indeed, defining the process efficiency by
$\eta\equiv\frac{E_{\rm extracted}}{E_{0}},$ (15)
we obtain efficiencies of $14.5\%$, $7.7\%$, and $1.2\%$ for the three cases $\hat{r}=1.2$, $1.5$, and $1.9$, respectively. The corresponding multi-particle approach by Piran reaches an efficiency of $12.6\%$ [21, 13, 12].
The main problem with the above results is that they apply only in the linear regime and neglect the role of the irreducible mass, which is mandatory as soon as the repetitive processes are considered (see next section).
IV A repetitive linear process
An alternative generalization of the Penrose process was suggested in Fig. 33.2 of [MTW73], moving from a single process to a repetitive one, owing to the scale invariance of the gravitational vacuum equations. Misner [as quoted in Ref. 7] had already applied this scale invariance in a different context. The authors attempt to achieve the program outlined by Penrose, who imagined “a civilization which has built some form of stabilized structure S surrounding the Black Hole” [see Fig. 5 in Ref. 7]. Indeed, they proposed, ” An advanced civilization has constructed a rigid framework around a black hole, and has built a huge city on that framework. Each day trucks carry one million tons of garbage out of the city to the garbage dump. At the dump the garbage is shoveled into shuttle vehicles which are then, one after another, dropped toward the center of the black hole”. Consequently, the process consists of repeating the decay by dropping many particles of the same kind, one by one, into the BH.
The ballistic process of Penrose is linear in the angular momentum and the mass, i.e., for a given initial state with mass $M_{i}$ and angular momentum $L_{i}$ the capture of a particle with energy $E_{1}$ and angular momentum $p_{\phi 1}$ leads to a new mass $M_{f}=M_{i}+E_{1}$ and new angular momentum $L_{f}=L_{i}+p_{\phi 1}$. Therefore, if one were to implement the advanced civilization idea by repeating $n$ times the capture process with the same particle with energy $E_{1}$, as reported in Figs. 13 and Table 1, one could naively assume that the final state of the BH would be $M_{f}(n)=M_{i}+nE_{1}$ and $L_{f}(n)=L_{i}+np_{\phi 1}$, appealing to a scale-invariant process. Then, one could raise the question about how many particles are necessary to extract the entire rotational energy of the BH, i.e., $L_{f}(n)=0$. For the case summarized in Table 1, the initial state is that of an extreme Kerr BH with $M_{i}=M$ and $L_{i}=M^{2}$, and the values of the captured particle are $E_{1}=-1.449\times 10^{-3}M$ and $p_{\phi 1}=-4.061\times 10^{-3}M$. Then, from the condition $L_{f}(n)=0$, one would obtain that $n\approx 246$, which corresponds to a final mass of $M_{f}(246)=0.643M$. Since the energy of the captured particle, $E_{1}$, is the result of the decay of an incoming particle with energy $E_{0}=10^{-2}M$, one would conclude that to reach the final state of a static Schwarzschild BH, it is necessary to inject $2.46M$ of energy! For the decays occurring at $\hat{r}=1.5$ and $\hat{r}=1.9$, the final state of a non-rotating BH is obtained after $236$ and $208$ decays, respectively. In all cases, this naive approach necessarily leads to a large violation of mass-energy conservation and is physically unacceptable. This paradoxical result is represented in Fig. 4 for the decay considered in Fig. 1 and Table 1.
These paradoxes prove the incompleteness of the above linear implementation of a repetitive Penrose ballistic process. To fulfill the mass-energy conservation, the solution must consider the nonlinear relation between the BH mass and angular momentum given by the mass-energy formula (see below Eq. 16), which is a function of the three independent parameters mass $M$, charge $e$, and angular momentum $L$ as well as of the irreducible mass $M_{\rm irr}$. Furthermore, the global consistency also needs the additional implementation of the law of the monotonic increase of the BH surface area given by Eq. (4) [8, 15, 16]. The reversible and irreversible transformations of a BH have to be considered. The concept of extractable energy, complementary to the extracted energy in the original Penrose process computed above, needs to be considered. Any decay process becomes dominated by nonlinear equations, and the scale invariance is broken.
Figure 4: A linear repetitive decay process occurring at $\hat{r}=1.2$ A number $n$ of particles with energy $E_{0}$ is injected into the BH. After $n\approx 246$ decays, the end state is a Schwarzschild BH of mass $\sim 0.643M$.
V Conclusions
We have reconsidered the basic constituents of the ballistic Penrose process and shown that it remains a remarkably simple method to probe the rotational energy extraction from a Kerr BH. We have first considered, for an extreme BH with $a/M=1$, the case of decay of a test particle of mass $\mu_{0}$ and initial angular momentum $p_{\phi 0}$ into two particles: $\mu_{1}$ with counter-rotating angular momentum and $\mu_{2}$ carrying away the extracted energy (see Sec. II). We perform this process at three different radii within the ergosphere, $\hat{r}=1.2$, $\hat{r}=1.5$, and $\hat{r}=1.9$, and we compare and contrast the results. So, we overcome the criticism addressed toward the original Penrose process. We confirm that it carries out the expected task of energy and angular momentum extraction in every single process with an efficiency of 14.5% (see Sec. III).
In Section. IV, we have considered the interesting proposal by [MTW73] to address a repetitive Penrose process at radii $\hat{r}=1.2$, $\hat{r}=1.5$, $\hat{r}=1.9$. Since the Penrose process is linear in the mass and angular momentum, one may consider performing such an iteration under the simplifying assumption that the energy extracted at each step coincides with the one at the first step evaluated for $a/M=1$. Under this assumption, one would obtain that the entire rotational energy can be extracted (see Fig. 4). However, energy conservation is violated. Indeed, at each step the extracted energy should be computed using the new value of $a/M$ derived by computing the new trajectories in the Kerr solution. This introduces a non-linearity in the original Penrose process, apparently linear in the mass $M$ and angular momentum $L$. This path is approached in the accompanying paper, where such nonlinearity is accounted for by considering that, in each step of the ballistic process, the extracted energy must be computed using the new value of $a/M$ derived from the new trajectories in the Kerr solution using the effective potential (1), instead of that given by Eq. (6) valid only for the extreme Kerr BH, and the BH mass-energy formula [15, 16]
$M^{2}=\left(M_{\rm irr}+\frac{e^{2}}{4M_{\rm irr}}\right)^{2}+\frac{L^{2}}{4M_{\rm irr}^{2}},$ (16)
where $M$ is expressed using a nonlinear equation involving the irreducible mass $M_{\rm irr}$ and the angular momentum $L$. The irreducible mass $M_{\rm irr}$ is directly related to the monotonically increasing surface area $S$ of the BH given by Eq. (4). There, the concepts of reversible and irreversible transformations are introduced. This leads to the proof of the irreversibility of the Penrose process. This will also need the introduction of a highly nonlinear extractable energy, in addition to the simple linear extracted energy discussed in the first part of this letter.
Crucial will be to compare and contrast in this new approach the above three sequences at $\hat{r}=1.2$, $\hat{r}=1.5$, $\hat{r}=1.9$. We use the Penrose process as a prototype to infer how the presence of the $M_{\rm irr}$ significantly affects the rotational energy extraction processes in pure self-gravitating systems.
Acknowledgements.
We thank Profs. C.L. Bianco, D. Christodoulou, R. Kerr, and J.A. Rueda for discussions on this work.